# Resource Allocation with Multi-Team Collaboration Based on Hamilton’s Rule

### v1

@@ I Introduction
The ability of multi-agent systems to collaborate effectively is critical for large-scale autonomous robotic systems, where dynamic and uncertain environments may require adaptive resource allocation across large mission areas. This distributed collaboration between multiple robot teams requires a careful balance between mission priorities for individual teams and overall system performance, e.g., [1, 2, 3, 4]. As an approach to solving this problem, we take inspiration from behaviors observed in nature, where altruistic decision-making between related organisms can increase the overall fitness of the group, so long as the benefit gained by the benefiting organism is greater than the relative cost to the altruistic organism [5]. In this sense, we consider how teams of robots may relate to each other in the context of relative mission importance, and we consider the use of individual robots themselves as a shared resource between all cooperating teams.
Resource allocation in multi-agent systems is a problem that has been approached in the literature to distribute computation resources [6, 7, 8], with applications including cyber-physical systems [9], UAV team deployment [10], and smart grid scheduling [11]. In addition to computation resource distribution, other works consider collaboration in manufacturing [12] and the collaborative management of supply chains [13]. However, in this work, we consider scenarios where agents themselves may be considered as a system resource, with the assignment of an agent to a team increasing the effectiveness of the team in completing their assigned mission.
In contrast with previous approaches, this work contributes to solving the resource allocation problem for multi-team systems by formulating the problem as an ecologically inspired altruistic decision process, where teams compare the relative importance of their assigned missions, as well as relative cost or benefit between teams gaining and losing robot team members, respectively. Thus, we propose a framework ensuring that resources are reallocated between teams only if the collaboration between them is beneficial for the entire multi-team system, and give an example of this framework applied to a multi-team coverage control problem.
@@
II Hamilton’s Rule for a Team-Level Collaboration Framework
A common inspiration for modeling systems is the environment and nature surrounding us; understanding how it moves and works has been an influence to previous literature work, as in [14, 15, 16, 17]. In this paper, we are inspired by and extract the properties of Hamilton’s rule [5]. Unlike prior work [18, 19] that focus on one team of cooperating agents, our problem setting in this paper considers multiple teams of agents, with each team initially assigned to operate in a distinct region and potentially allocating robots to different teams for the benefit of the entire multi-team system. To that end, graph theory can provide an appropriate framework for modeling and analyzing interactions among those teams [20]. Let $G=(V,E)$ be the undirected graph abstracting the $m$ different available teams of agents as well as their potential interactions, where $V=\{v_{1},v_{2},\dots,v_{m}\}$ is the set of nodes representing the teams, and $E\subseteq V\times V$ is the set of edges representing the potential collaborative interactions between them.
@@ II-A Pairwise Uni-Directionality of Hamilton’s Rule
where $F_{k}(n_{k}+1)$ denotes the value of team $k$’s evaluation function after receiving an agent, and $F_{k}(n_{k}-1)$ denotes the value after losing an agent. Since $F_{k}$ is strictly increasing, we have:
@@
We prove the result by contradiction. Assume that both
$r_{ij}\,B_{j}>C_{i}\quad\text{and}\quad r_{ji}\,B_{i}>C_{j}$
hold simultaneously. Writing these inequalities in terms of weights, we get:
$w_{j}\,B_{j}>w_{i}\,C_{i}$ (5)
$w_{i}\,B_{i}>w_{j}\,C_{j}$ (6)
From (5), we deduce that:
$w_{j}>w_{i}\,\frac{C_{i}}{B_{j}}$ (7)
Substituting (7) into (6), we obtain:
@@
Canceling $w_{i}$ (with $w_{i}>0$), yields:
@@
by (2) and (3). Thus, we have:
$B_{i}\leq C_{i}\quad\text{and}\quad B_{j}\leq C_{j}.$ (9)
We can rewrite the inequalities in (9) into the following equalities: $C_{i}=B_{i}+\alpha$ and $C_{j}=B_{j}+\beta$, such that $\alpha,\beta\geq 0$. Then,
$C_{i}\,C_{j}$ $=(B_{i}+\alpha)(B_{j}+\beta)$
$=B_{i}\,B_{j}+\beta B_{i}+\alpha B_{j}+\alpha\beta.$
@@
which contradicts inequality (8). Thus, since all agents are identical, i.e., have equal “resource contribution”, under Assumption 1, the assumption that both inequalities hold must be false, proving Theorem 1. ∎
@@ II-B Multi-Team Bidding Process
In Section II-A, we focused on pairwise collaborations using Hamilton’s Rule. However, when considering a system with multiple teams, the problem is inherently more challenging. Instead of only evaluating local, pairwise interactions, the overall system performance must be taken into account. We say that given two teams $i,j\in E$, an outgoing collaboration from team $i$ to team $j$ and an incoming collaboration to team $i$ from team $j$ is beneficial when Equation 1 holds. We must note that if a team ends up with 00 robots after collaborating, in the next collaboration process, that team will only be considered for incoming collaborations.
@@ II-C Iterative Collaboration Framework
$\mathcal{G}(n_{1}^{*},\dots,n_{m}^{*})\;\;\geq\;\;\mathcal{G}(n_{1},\dots,n_{m})$
$\forall\,(n_{1},\dots,n_{m})\in\Omega.$
@@
At each iteration $t$, if at least one collaboration (transfer) is executed, then by the design of Algorithm 1 we have:
@@
Since $\Omega$ is finite and $\mathcal{G}$ is a real-valued function on $\Omega$, by Lemma 1 the function $\mathcal{G}$ attains a maximum value on $\Omega$, say
@@
Thus, the strictly increasing sequence $\{\mathcal{G}(x^{(t)})\}$ is bounded above by $M$. Because the sequence takes values in a finite set, it cannot increase indefinitely and must eventually stabilize; that is, there exists some $T\in\mathbb{N}$ such that:
$\mathcal{G}\bigl{(}x^{(T+1)}\bigr{)}\leq\mathcal{G}\bigl{(}x^{(T)}\bigr{)}.$
Since a collaboration is executed only when it yields a strict increase in $\mathcal{G}$ (see Line 22 of Algorithm 1), we have that at iteration $T$ no allowed transfer can produce a strict increase. Thus, $x^{(T)}$ is a local optimum with respect to the allowed moves. Next, we show that under Assumption 2, any local optimum with respect to these transfers is in fact a global optimum. Suppose for contradiction that $x^{(T)}$ is not a global optimum. Then there exists some allocation $x^{*}\in\Omega$ such that
@@
Since the global objective $\mathcal{G}$ is point-wise concave in each coordinate (owing to the point-wise concavity of each $F_{k}$ by Assumption 2 and the fact that $w_{k}>0$), the function $\mathcal{G}$ has the property that any local improvement —i.e., any feasible agent transfer from one team to another that is allowed by Hamilton’s Rule— would yield a strictly higher value of $\mathcal{G}$. Therefore, the existence of an allocation $x^{*}$ with $\mathcal{G}(x^{*})>\mathcal{G}(x^{(T)})$ implies that there exists at least one allowed transfer from $x^{(T)}$ that would increase $\mathcal{G}$, contradicting the termination condition of the algorithm. Thus, the terminal allocation $x^{(T)}$ must coincide with a global maximizer of $\mathcal{G}$ on $\Omega$. Since the process only permits transfers that strictly increase $\mathcal{G}$ and $\Omega$ is finite, the iterative process converges to the optimal allocation in a finite number of steps, thus proving our theorem. ∎
@@ III Team Mission: Area Coverage
To illustrate the operation of the proposed framework, we consider the mission assigned to each robot team to be optimal area coverage, to which a typical approach is the Voronoi-based coverage control introduced in [22]. Specifically, we consider the deployment of a team of $N\in\mathbb{Z}^{+}$ robots to cover a convex and compact domain of interest $\mathcal{D}\subset\mathbb{R}^{d}$, where $d\in\mathbb{Z}^{+}$, associated with a density function $\phi(q):\mathcal{D}\rightarrow\mathbb{R}^{+}$ encoding the relative importance of different points in $\mathcal{D}$. By furthermore assuming the farther a robot is from a point $q\in\mathcal{D}$, the less effectively it covers $q$, the locational cost function can be formulated as [22]
[…]
@@
$\mathcal{V}_{i}$ =$=\left\{{q}\in\mathcal{D}\,|\,\lVert{q}-p_{i}\rVert\leq\lVert{q}-p_{j}\rVert,\,\,\forall j\neq i\in\mathcal{N}\right\}$ (17)
@@
A typical approach to find an optimal coverage configuration of the robots, e.g., [22], is that robot $i$ follows the gradient flow of $L(p)$ with respect to $p_{i}$, $\forall i\in\mathcal{N}$, which is known as a continuous-time version of Lloyd’s algorithm, asymptotically leading to a centroidal Voronoi tessellation (CVT), i.e.,
[…]
@@
[…]
To meet Assumption 2, we let the mission evaluation function $F(N)=-L(N,c)$ for some $c$. In Theorem 3 and Corollary 19, we investigate the strictly decreasing property of $L(N,c)$ with respect to $N$, i.e., the strictly increasing property of $F(N)$ presented in Assumption 2.
@@
Suppose $p_{i}\in\mathcal{D}$, $\forall i\in\mathcal{N}$. By adding one more robot $p_{N+1}\in\mathcal{D}\setminus\{p_{i}\}$, $\forall i\in\mathcal{N}$, the locational cost (18) decreases, i.e.,
@@
Based on [22], the locational cost function (18) can be rewritten as,
[…]
By adding one more robot, as illustrated in Fig.3, since $\phi(q)\in\mathbb{R}^{+}$, $\forall q\in\mathcal{D}$, and
$\min_{i\in\mathcal{N}^{\prime}}\|p_{i}-q\|\leq\min_{i\in\mathcal{N}}\|p_{i}-q\|,$
where $\mathcal{N}^{\prime}=\mathcal{N}\cup\{N+1\}$, then we have
$L(N,\,p)\geq L(N+1,\,[p^{\top},\,p^{\top}_{N+1}]^{\top}).$
Furthermore, since $p_{N+1}\in\mathcal{D}\setminus\{p_{i}\}$, $\forall i\in\mathcal{N}$, then $\exists\,\mathcal{A}\subset\mathcal{D}$ such that the Lebesgue measure $\lambda(\mathcal{A})\neq 0$, $p_{N+1}\in\mathcal{A}$, and
$\|p_{N+1}-q\|<\|p_{i}-q\|,\quad\forall q\in\mathcal{A},\,\forall i\in\mathcal{N},$
then we have
$L(N,\,p)>L(N+1,\,[p^{\top},p^{\top}_{N+1}]^{\top}),\quad\forall N\in\mathbb{Z}^{+},$
proving Theorem 3. ∎
Figure 4: The locational cost (18) at a CVT with $\phi(q)=e^{-(\frac{q_{x}^{2}}{0.8^{2}}+\frac{q_{y}^{2}}{0.8^{2}})}$ with respect to the total number of robots $N$.
@@
Suppose $N$ robots have converged to a CVT, i.e., $p_{i}=c_{i}$, $\forall i\in\mathcal{N}$. After adding one more robot $p_{N+1}\in\mathcal{D}\setminus\{p_{i}\}$, $\forall i\in\mathcal{N}$, a new CVT $c^{\prime}$ with $N+1$ robots resulting from the Lloyd’s algorithm has a lower locational cost value than the old CVT $c$ with $N$ robots does, i.e.,
$L(N+1,c^{\prime})<L(N,c),\quad\forall N\in\mathbb{Z}^{+}.$ (19)
@@
Suppose the $N\in\mathbb{Z}^{+}$ robots have converged to a CVT $c$, by adding $p_{N+1}\in\mathcal{D}\setminus\{p_{i}\}$, $\forall i\in\mathcal{N}$, according to Theorem 3, we have
$L(N,c)>L(N+1,[c^{T},p^{T}_{N+1}]^{T}).$ (20)
Then, Lloyd’s algorithm, e.g., [22], results in
$L(N+1,[c^{T},p^{T}_{N+1}]^{T})\geq L(N+1,c^{\prime}),$ (21)
@@
Figure 5: Simulation of Algorithm 1 with $m=4$ and $N=16$ while having all teams have equal weights, but different mission evaluation functions by changing the density function $\phi(q)$ in the locational cost in Equation 16. (a) shows the initial random allocation of robots to teams, (b) shows the final robot allocation, and (c) shows the value of $\mathcal{G}$ at each collaboration iteration until collaboration is no longer possible (hence, getting the allocation in (b)).
The property of diminishing returns of the mission evaluation function $F(N)=-L(N,c)$ is not trivial to prove in the context of Voronoi-based coverage control. Nevertheless, we empirically find that $L(N,c)$ asymptotically decreases with respect to $N$ with a diminishing rate, as illustrated in Fig.4, which corresponds to the property of diminishing returns of $F(N)$ presented in Assumption 2. We leave its theoretical proof as a future work.

### v2

@@ I Introduction
Multi-robot systems are increasingly deployed in large-scale and safety-critical applications such as search-and-rescue, environmental monitoring, and disaster response, e.g., [1, 2, 3, 4, 5, 6]. Some of these tasks often require teams of robots to operate in parallel across geographically distinct regions, while balancing between mission priorities for individual teams and overall system performance. A key challenge in such scenarios is the optimal allocation of available robots across multiple teams, weighed by the priorities of the individual team missions, driving the entire system to an optimized overall effectiveness.
This challenge connects directly to the broader literature on resource allocation in multi-agent systems, e.g., [7, 8, 9], with applications including cyber-physical systems [10], UAV team deployment [11], and smart grid scheduling [12]. For instance, [11] combines reinforcement learning with communication resource allocation, for UAV deployment, optimizing coverage and communication. In addition to computation resource distribution, other works consider collaboration in manufacturing [13] and the collaborative management of supply chains [14]. Game-theoretic frameworks have also been proposed to handle dynamically changing environments, where task demands evolve over time and robot allocations must adapt accordingly, e.g., [15]. In practice, however, assigning an additional robot to a team can directly increase that team’s mission effectiveness.
A common inspiration for modeling systems is the environment and nature surrounding us; understanding how it moves and works has been an influence to previous literature work, as in [16, 17, 18, 19]. In this paper, we are inspired by and extract the properties of Hamilton’s rule [20]. This rule, originally formulated in evolutionary biology, states that altruistic behavior emerges when the weighted benefit to others exceeds the cost to the donor. By analogy, we consider how teams of robots may relate to each other in terms of relative mission importance, and we view the robots themselves as transferable resources that may be shared among cooperating teams.
Building on this perspective, we design a framework in which teams compare both the relative importance of their missions and the marginal cost or benefit of reallocating robots. Transfers are permitted only when they improve the performance of the overall multi-team system. We demonstrate this idea in the context of a Voronoi-based coverage control problem, e.g., [21], showing how altruistic decisions lead to improved system outcomes. Unlike prior work [22] that focused on a single team of cooperating robots, our setting explicitly addresses multiple teams of robots, each initially assigned to a distinct region but capable of reallocating robots for a collective benefit.
@@
II A Team-Level Collaboration Framework Based on Hamilton’s Rule
In this section, we formulate the problem of resource allocation in multi-agent teams based on Hamilton’s rule. To begin with, we introduce a graph-based model, e.g., [23], to represent team interactions. Let $G=(V,E)$ be the undirected graph abstracting the $m$ different available teams of agents as well as their potential interactions, where $V=\{v_{1},v_{2},\dots,v_{m}\}$ is the set of nodes representing the teams, and $E\subseteq V\times V$ is the set of edges representing the potential collaborative interactions between them.
@@ II-A Pairwise Uni-Directionality of Hamilton’s Rule
where $F_{k}(n_{k}+1)$ denotes the value of team $k$’s evaluation function after receiving an agent, and $F_{k}(n_{k}-1)$ denotes it after losing an agent. Since $F_{k}$ is strictly increasing, we have:
@@
We prove the result by contradiction. Assume that both $r_{ij}\,B_{j}>C_{i}$ and $r_{ji}\,B_{i}>C_{j}$ hold simultaneously. Writing these two inequalities in terms of weights, we get $w_{j}\,B_{j}>w_{i}\,C_{i}$ and $w_{i}\,B_{i}>w_{j}\,C_{j}$, leading to
@@
Canceling $w_{i}$ (with $w_{i}>0$) yields
@@
by (2) and (3). Thus, we have $B_{i}\leq C_{i}$ and $B_{j}\leq C_{j}$, which can be rewritten as $C_{i}=B_{i}+\alpha$ and $C_{j}=B_{j}+\beta$, such that $\alpha,\beta\geq 0$. Then,
$C_{i}\,C_{j}=(B_{i}+\alpha)(B_{j}+\beta)=B_{i}\,B_{j}+\beta B_{i}+\alpha B_{j}+\alpha\beta.$
@@
which contradicts the inequality (5). Thus, since all agents are identical, i.e., have equal “resource contribution”, under Assumption 1, the assumption that both inequalities hold must be false, proving Theorem 1. ∎
@@ II-B Multi-Team Bidding Process
In Section II-A, we focused on pairwise collaborations using Hamilton’s Rule. However, when considering a system with multiple teams, the problem is inherently more challenging. Instead of only evaluating local, pairwise interactions, the overall system performance must be taken into account. We say that given two teams $i,j\in E$, an outgoing collaboration from team $i$ to team $j$ and an incoming collaboration to team $i$ from team $j$ is beneficial when Equation 1 holds. We must note that if a team ends up with 0 robots after collaborating, in the next collaboration process, that team will only be considered for incoming collaborations.
@@ II-C Iterative Collaboration Framework
$\mathcal{G}(n_{1}^{*},\dots,n_{m}^{*})\;\;\geq\;\;\mathcal{G}(n_{1},\dots,n_{m}),\quad\forall\,(n_{1},\dots,n_{m})\in\Omega.$
@@
At each iteration $t$, if at least one collaboration (transfer) is executed, then by the design of Algorithm 1, we have
@@
Since $\Omega$ is finite and $\mathcal{G}$ is a real-valued function on $\Omega$, by Lemma 1 the function $\mathcal{G}$ attains a maximum value on $\Omega$, denoted by
@@
Thus, the strictly increasing sequence $\{\mathcal{G}(x^{(t)})\}$ is bounded above by $M$. Because the sequence takes values in a finite set, it cannot increase indefinitely and must eventually stabilize; hence, no further collaborations would occur from that point on. More formally, there exists some $T\in\mathbb{N}$ such that
$\mathcal{G}\bigl{(}x^{(T+1)}\bigr{)}\leq\mathcal{G}\bigl{(}x^{(T)}\bigr{)},$
where no further collaborations occur at $T+1$. Since a collaboration is executed only when it yields a strict increase in $\mathcal{G}$ (see Line 22 of Algorithm 1), we have that at iteration $T$ no allowed transfer can produce a strict increase. Thus, $x^{(T)}$ is a local optimum with respect to the allowed moves.
Next, we show that under Assumption 2, any local optimum with respect to these transfers is in fact a global optimum. Suppose for contradiction that $x^{(T)}$ is not a global optimum. Then there exists some allocation $x^{*}\in\Omega$ such that
@@
Since the global objective $\mathcal{G}$ is point-wise concave in each coordinate (owing to the point-wise concavity of each $F_{k}$ by Assumption 2 and the fact that $w_{k}>0$), the function $\mathcal{G}$ has the property that any local improvement, i.e., any feasible agent transfer from one team to another that is allowed by Hamilton’s Rule, would yield a strictly higher value of $\mathcal{G}$. Therefore, the existence of an allocation $x^{*}$ with $\mathcal{G}(x^{*})>\mathcal{G}(x^{(T)})$ implies that there exists at least one allowed transfer from $x^{(T)}$ that would increase $\mathcal{G}$, contradicting the termination condition of the algorithm. Thus, the terminal allocation $x^{(T)}$ must coincide with a global maximizer of $\mathcal{G}$ on $\Omega$. Since the process only permits transfers that strictly increase $\mathcal{G}$ and $\Omega$ is finite, the iterative process converges to the optimal allocation in a finite number of steps, thus proving our theorem. ∎
@@ III Team Mission: Area Coverage
To illustrate the operation of the proposed framework, we consider the mission assigned to each robot team to be optimal area coverage, to which a typical approach is the Voronoi-based coverage control introduced in [21]. Specifically, we consider the deployment of a team of $N\in\mathbb{Z}^{+}$ robots to cover a domain of interest $\mathcal{D}\subset\mathbb{R}^{d}$, which is a measurable set with nonempty interior, i.e., $\mathrm{Int}(\mathcal{D})\neq\emptyset$, associated with a measurable density function $\phi(q):\mathcal{D}\rightarrow\mathbb{R}^{+}$ encoding the relative importance of different points in $\mathcal{D}$. By assuming the farther (based on the Euclidean norm) a robot is from a point $q\in\mathcal{D}$, the less effectively it covers $q$, the locational cost function can be formulated as [21]
$L(p)=\sum_{i\in\mathcal{N}}\int_{\mathcal{V}_{i}}\!\|p_{i}-q\|^{2}\phi(q)\,dq,$ (12)
@@
$\mathcal{V}_{i}$ $=\left\{{q}\in\mathcal{D}\,|\,\lVert{q}-p_{i}\rVert\leq\lVert{q}-p_{j}\rVert,\,\,\forall j\neq i\in\mathcal{N}\right\}$ (13)
@@
A typical approach to find an optimal coverage configuration of the robots, e.g., [21], is that robot $i$ follows against the gradient of $L(p)$ with respect to $p_{i}$, $\forall i\in\mathcal{N}$, which is known as a continuous-time version of Lloyd’s algorithm, asymptotically leading to a centroidal Voronoi tessellation (CVT), i.e.,
$p_{i}=c_{i}=\frac{\int_{\mathcal{V}_{i}}\!{q}\phi({q})\,d{q}}{\int_{\mathcal{V}_{i}}\!\phi({q})\,d{q}},\quad\forall i\in\mathcal{N},$
@@
$L(N,p)=\sum_{i\in\mathcal{N}}\int_{\mathcal{V}_{i}}\!\|p_{i}-q\|^{2}\phi(q)\,dq.$ (14)
In addition, Fig.3 illustrates an example of how the Voronoi tessellation changes after adding one robot to a team of robots.
To meet Assumption 2, we let the mission evaluation function F(N)=L(N,c)F(N)=-L(N,c) for some $c$. In Theorem 3 and Corollary 1, we investigate the strictly decreasing property of $L(N,c)$ with respect to $N$, i.e., the strictly increasing property of $F(N)$ presented in Assumption 2.
@@
Suppose $p_{i}\in\mathcal{D}$, $\forall i\in\mathcal{N}$. By adding one more robot $p_{N+1}\in{\mathrm{Int}(\mathcal{D})}\setminus\{p_{i}\}$, $\forall i\in\mathcal{N}$, the locational cost (14) decreases, i.e.,
@@
Based on [21], the locational cost function (14) can be rewritten as
$L(N,p)=\int_{\mathcal{D}}\min_{i\in\mathcal{N}}\|p_{i}-q\|^{2}\phi(q)\,dq.$
In addition, we denote
$H(p^{\prime},q)=\min_{i\in\mathcal{N}}\|p_{i}-q\|^{2}-\min_{i\in\mathcal{N}^{\prime}}\|p_{i}-q\|^{2},$
where $\mathcal{N}^{\prime}=\mathcal{N}\cup\{N+1\}$. Apparently, $H(p^{\prime},q)\geq 0$, $\forall q\in\mathcal{D}$.
Moreover, we let
$\bar{\delta}=\min_{i\in\mathcal{N}}\|p_{N+1}-p_{i}\|>0.$
Since $p_{N+1}\in\mathrm{Int}(\mathcal{D})$, then $\exists\,{\varepsilon}>0$ such that the open Euclidean ball
$\mathcal{B}(p_{N+1},{\varepsilon})=\{q\in\mathbb{R}^{d}\mid\|q-p_{N+1}\|<\varepsilon\}\subset\mathcal{D}.$
Furthermore, we define
$\bar{\varepsilon}=\min\{\varepsilon,\bar{\delta}/2\}\quad\text{and}\quad\bar{\mathcal{B}}=\mathcal{B}(p_{N+1},\bar{\varepsilon}).$
Then, the Lebesgue measure $\lambda(\bar{\mathcal{B}})>0$, and thus we have
$\|q-p_{i}\|\geq$ $\,\|p_{i}-p_{N+1}\|-\|q-p_{N+1}\|$
$>$ $\,\bar{\delta}-\bar{\varepsilon}$
$\geq$ $\,\bar{\delta}/2>\|q-p_{N+1}\|,\quad\forall q\in\bar{\mathcal{B}},\,\forall i\in\mathcal{N}.$
Hence, we get
$\min_{i\in\mathcal{N}^{\prime}}\|p_{i}-q\|^{2}=\|q-p_{N+1}\|^{2}<\min_{i\in\mathcal{N}}\|p_{i}-q\|^{2},\quad\forall q\in\bar{\mathcal{B}},$
which means that $H(p^{\prime},q)>0$, $\forall q\in\bar{\mathcal{B}}$. Together with the assumption $\phi(q)>0$ and the fact $H(p^{\prime},q)\geq 0$, $\forall q\in\mathcal{D}$, we can have
$L(N,p)-L(N+1,p^{\prime})$
$=$ $\int_{\mathcal{D}}\!H(p^{\prime},q)\phi(q)\,dq$
$=$ $\int_{\bar{\mathcal{B}}}\!H(p^{\prime},q)\phi(q)\,dq+\int_{\mathcal{D}\setminus\bar{\mathcal{B}}}\!H(p^{\prime},q)\phi(q)\,dq$
$\geq$ $\int_{\bar{\mathcal{B}}}\!H(p^{\prime},q)\phi(q)\,dq>0,$
which proves $L(N+1,p^{\prime})<L(N,p)$. ∎
Figure 4: The locational cost (14) at a CVT with $\phi(q)=e^{-(\frac{q_{x}^{2}}{0.8^{2}}+\frac{q_{y}^{2}}{0.8^{2}})}$ with respect to the total number of robots NN.
@@
Suppose $N$ robots have converged to a CVT, i.e., $p_{i}=c_{i}$, $\forall i\in\mathcal{N}$. After adding one more robot $p_{N+1}\in{\mathrm{Int}(\mathcal{D})}\setminus\{p_{i}\}$, $\forall i\in\mathcal{N}$, a new CVT $c^{\prime}$ with $N+1$ robots resulting from the Lloyd’s algorithm has a lower locational cost value than the old CVT $c$ with $N$ robots does, i.e.,
$L(N+1,c^{\prime})<L(N,c),\quad\forall N\in\mathbb{Z}^{+}.$
@@
Suppose the $N\in\mathbb{Z}^{+}$ robots have converged to a CVT $c$, by adding $p_{N+1}\in{\mathrm{Int}(\mathcal{D})}\setminus\{p_{i}\}$, $\forall i\in\mathcal{N}$, as per Theorem 3, we have $L(N,c)>L(N+1,[c^{T},p^{T}_{N+1}]^{T})$. Then, the Lloyd’s algorithm, e.g., [21], results in
$L(N+1,[c^{T},p^{T}_{N+1}]^{T})\geq L(N+1,c^{\prime}),$
@@
Figure 5: Simulation of Algorithm 1 with $m=4$ and $N=16$ while having all teams have equal weights, but different mission evaluation functions by changing the density function $\phi(q)$ in the locational cost in Equation 12. (a) The initial random allocation of robots to teams. (b) The final robot allocation. (c) The value of $\mathcal{G}$ at each collaboration iteration until collaboration is no longer possible (hence, getting the allocation in (b)).
The property of diminishing returns of the mission evaluation function F(N)=L(N,c)F(N)=-L(N,c) is not trivial to prove in the context of Voronoi-based coverage control. Nevertheless, we empirically find that $L(N,c)$ asymptotically decreases with respect to $N$ with a diminishing rate, as illustrated in Fig.4, which corresponds to the property of diminishing returns of $F(N)$ presented in Assumption 2. We leave its theoretical proof as a future work.